OpenStudy (nerdgamer_):

James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done? three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths

OpenStudy (nerdgamer_):

\(\Huge Please~Help!\)

OpenStudy (solomonzelman):

Well, when you add fractions, you add the whole numbers together, and then add the fractions. Are you familiar with adding fractions?

OpenStudy (nerdgamer_):

sort of :3

OpenStudy (nerdgamer_):

\(\Large I~NEED~HELP!!!!!!!!\)

OpenStudy (zubhanwc3):

\[3\frac{ 2 }{ 3 } + 4\frac{ 1 }{ 8 } = 7\frac{ 3 }{ 11 }\] it is wrong, he didnt get common a denominator so the asnwer should be \[7(\frac{ 16 }{ 24 }+\frac{ 3 }{ 24 }) = 7\frac{ 19 }{ 24 }\]

OpenStudy (zubhanwc3):

just came on, so srry for the late reply